Please help me solve this problem:
I need to solve the formula below for the indicated variable.
E
_______________ = I, for R
R + r

Answer by checkley77(12844)

E
_______________ = I, for R
R + r
I(R+r)=E
IR+Ir=E
IR=E-Ir
R=(E-Ir)/I ANS.